# If Probable, Then Acceptable? Understanding Conditional Acceptability Judgments in Large Language Models

### v1

@@ Abstract
Jasmin Orth1   Philipp Mondorf1,2   Barbara Plank1,2 1MaiNLP, Center for Information and Language Processing, LMU Munich, Germany 2Munich Center for Machine Learning (MCML), Munich, Germany jasmin.orth@campus.lmu.de, {p.mondorf, b.plank}@lmu.de
@@ 1 Introduction
While both sentences are grammatically sound, the second appears less acceptable: it is unlikely that Mark wearing socks has any effect on whether his TV is working. This notion of acceptability, or how appropriate a conditional statement seems in a given context, is often judged intuitively by humans (Douven and Verbrugge, 2012; Skovgaard-Olsen et al., 2016). Two main hypotheses aim to explain such intuitions. The conditional probability hypothesis holds that acceptability depends primarily on the conditional probability of $B$ given $A$ (Evans et al., 2003; Over et al., 2007). By this account, both conditionals above might be judged acceptable, as the probability of the TV being turned off given that the plug has been pulled, and the probability of the TV working given that Mark is wearing socks, are both high. In contrast, the evidentiality hypothesis emphasizes an additional influence of relevance or evidential connection encoded within the statement (Douven and Verbrugge, 2012; Skovgaard-Olsen et al., 2016). From this perspective, (i) is acceptable because pulling the plug is causally connected to the TV turning off, whereas (ii) is not, because wearing socks lacks such a connection. Large language models have demonstrated notable capabilities across diverse linguistic and cognitive tasks, including general reasoning (Hao et al., 2024; Mondorf and Plank, 2024a), linguistic inference (Clark et al., 2020), and causal reasoning (Dettki et al., 2025; Yu et al., 2024). However, despite increasing work on LLM reasoning, research on conditional acceptability judgments in LLMs remains virtually nonexistent. Most existing studies on LLMs and conditional statements focus on their causal, abductive, or inferential capabilities (Liu et al., 2023; Dettki et al., 2025; Holliday et al., 2024; Mondorf and Plank, 2024b), rather than on whether the conditionals are judged contextually acceptable.
We address this gap by conducting a targeted investigation of conditional acceptability judgments in LLMs. Building on prior work from cognitive psychology (Skovgaard-Olsen et al., 2016), we assess the acceptability judgments of two model families: Llama 3.1 (8B and 70B) and Qwen 2.5 (7B and 72B), and examine how their judgments vary as a function of both the conditional probability of a statement and the evidential connection encoded within it. Our findings reveal that:
  • The acceptability judgments of models are influenced by both conditional probability and semantic relation, though the strength and consistency of this effect vary substantially by model family, size, and prompting technique.
@@ 2.1 Conditional Acceptability
The key concept of this study is conditional acceptability: the degree to which a conditional statement is perceived to be appropriate, plausible, or natural in context (Skovgaard-Olsen et al., 2016). While often associated with conditional probability $P(B\mid A)$, such as proposed by the conditional probability hypothesis (Evans et al., 2003; Oaksford and Chater, 2003; Over et al., 2007), the two can diverge significantly in practice. Take, for instance, the statement from the introduction: If Mark is wearing socks ($A$), then his TV will be working ($B$). This may be judged unacceptable or odd, even when both $P(A)$ and $P(B)$ are high, and even when $P(B\mid A)$ is high. The issue lies not in probability, but in the perceived irrelevance of $A$ to $B$. The statement violates intuitive notions of coherence or causality: wearing socks does not cause the TV to function, nor would not wearing them affect the outcome. As such, the distinction reflects the broader gap between correlation and causation (Barrowman, 2014). This disconnect has prompted extensive psychological research on how humans evaluate conditional statements (e.g., Douven, 2015; Evans and Over, 2004; Hardman and Macchi, 2003), suggesting that individuals often rely on a combination of semantic, contextual, and probabilistic cues when assessing conditionals (Berto and Özgün, 2021; Douven and Verbrugge, 2012; Over et al., 2007). Specifically, the evidentiality hypothesis argues that the semantic relation between antecedent $A$ and consequent BB—whether AA supports, contradicts, or is irrelevant to BB—plays a crucial role in acceptability judgments (Douven and Verbrugge, 2012; Skovgaard-Olsen et al., 2016; Berto and Özgün, 2021). Returning to the earlier example, If Mark is wearing socks, then his TV will be working may be judged unacceptable not because $P(B\mid A)$ is low, but because there is no plausible explanatory or causal (supporting) link between the two clauses. In particular, Skovgaard-Olsen et al. (2016) find that humans generally judge conditionals with a supporting relation more acceptable than those with a contradicting or irrelevant relation.
@@ 2.2 Dataset
To investigate how LLMs assess the acceptability of conditional statements, we draw on a dataset from  Skovgaard-Olsen et al. (2016). Their study aims to understand how humans judge conditional acceptability using 144 conditional statements embedded in 12 short, realistic contexts. [^1] More information about the dataset and demographics can be found in Appendix A.1 Each context is tied to a unique everyday scenario, such as going on a date or attending a meeting.
To enable systematic analyses, the 12 conditionals that share the same context are organized along two dimensions. First, each conditional can be classified into one of three relation types, indicating the semantic connection between antecedent $A$ and consequent $B$: a positive relation ($POS$) means $A$ supports $B$, a negative relation ($NEG$) denotes that $A$ undermines or contradicts $B$, and an irrelevant relation ($IRR$) means $A$ is unrelated to $B$. Second, the conditionals vary in terms of prior probability: each scenario includes four levels representing combinations of high or low probability for both antecedent $P(A)$ and consequent $P(B)$ (e.g., high-high, high-low, low-high, low-low; see Appendix A.1 for more details). This yields 12 unique configurations per context (3 relation types $\times$ 4 prior probabilities, as shown in Table 3 in the Appendix), resulting in 144 conditionals in total. The following example illustrates how relation types can be defined within a given context:
@@
As shown in Figure 2, a positive relation is illustrated by the conditional statement: If Nicole sunbathes on the beach, then she will get sunburned. Here, under the contextual constraint of her having sensitive skin, Nicole sunbathing actively causes her to get sunburned. In the negative example, the antecedent (given the context) contradicts her staying pale, while the irrelevant example pairs two unrelated ideas: knowing what a computer is has no causal connection to getting sunburned. These structured scenarios are designed to help participants apply real-world knowledge while simultaneously constraining interpretation, anchoring relations and probability judgments to a shared context. In the present study, we use this dataset to evaluate how LLMs judge the acceptability of conditionals and whether they exhibit similar patterns as humans. By prompting models with these same context-rich scenarios and conditionals, we can systematically assess their responses across relation types and probability levels.
@@ 3.1 Experimental Setup
We follow the approach of Skovgaard-Olsen et al. (2016) and present the 144 conditional statements described in Section 2.2 to LLMs in three distinct judgment tasks: (i) estimating the conditional probability $P(B\mid A)$ of the consequent given the antecedent, (ii.a) predicting the overall probability of the conditional $P\left(\text{``\emph{If A, then B}''}\right)$, and (ii.b) judging the acceptability of the conditional statement “If A, then B”. While tasks (ii.a) and (ii.b) present the LLMs with full conditional statements (e.g., ”If Mark is wearing socks, then his TV will be working”), the conditional probability task (i) presents the statement in a split format, consisting of a supposition such as “Suppose Mark is wearing socks.” and an outcome, such as ”His TV will be working”, which needs to be judged (see Figure 3 lower bounding box).
To analyze whether LLMs’ acceptability judgments align with a probabilistic or relevance-based interpretation, these tasks are designed to separate probabilistic reasoning from pragmatic or communicative considerations. The conditional probability task described above provides a reference point reflecting pure probabilistic reasoning: if large language models rely solely on conditional probability when judging conditional statements, their if-probability [^2] To make the distinction between the conditional probability $P(B\mid A)$ and the probability of the full conditional statement as clear as possible, we denote the probability of the full conditional statement the if-probability in this study. and acceptability ratings would be expected to closely match their corresponding $P(B\mid A)$ values. The split between acceptability and if-probability judgments further allows to determine whether assessments of conditional statements differ depending on the type of judgment. All tasks instruct the model to respond with a single number between 01000-100, indicating the degree of probability or acceptability. For a discussion on extracting probability judgments through direct elicitation as opposed to other techniques such as sentence probability, please refer to Appendix A.4. Full prompt templates are provided in Figures 8 to 10 in Appendix A.3.
Post-Processing
@@ 3.2 Prompts & Models
We aim to examine the impact of model family, model size, and prompting technique on the conditional acceptability judgments of LLMs. Therefore, we assess a total of four language models: Llama-3.1-8B-Instruct, [^3] https://huggingface.co/meta-llama/models Llama-3.1-70B- Instruct,footnotemark: Qwen2.5-7B-Instruct, [^4] https://huggingface.co/Qwen/collections and Qwen2.5-72B-Instruct,footnotemark: with the publicly accessible weights obtained from the Hugging Face platform. All models are prompted using a vanilla (zero-shot) and three-shot format, while the two larger models (Llama 70B and Qwen 72B) are also tested with chain-of-thought (CoT) prompting. To account for potential outliers, all models are prompted 5 times per data sample. An analysis of the consistency of model responses can be found in Appendix A.5. For further details on models and prompts, please refer to Appendix A.2.
Figure 4: Distribution of center-scaled judgments across humans and LLMs. Top row: overall distributions. Bottom row: distributions divided by relation types. Left: Llama 70B (vanilla), center: Qwen 72B (vanilla), right: humans.
@@ 3.3 Statistical Methods
Mixed-effects models allow for the analysis of how conditional probability and relation type influence conditional acceptability judgments while accounting for additional sources of variation. In particular, different scenarios, participants, [^5] While we do not collect human data ourselves, we use human data from Skovgaard-Olsen et al. (2016) to compare LLM trends. and LLM sampling iterations introduce underlying variation that might otherwise obscure meaningful effects. Linear mixed-effects models account for this structure by incorporating both fixed effects (the predictors of interest) and random effects (uncontrolled variation across groups) (Brown, 2021). Further information on model fitting can be found in Appendix A.6.
@@ 4 Results
Impact of Conditional Probability
For all LLMs, judgments of conditional statements are strongly influenced by the underlying conditional probability $P(B\mid A)$. ANOVA results (Tables 10 to 18 in the Appendix) confirm a highly significant main effect of conditional probability across all models ($p<0.0001$). As shown in Figure 6, judgments increase with rising conditional probability, a pattern further supported by fixed effects estimates in Table 1, which quantify the effect of $P(B\mid A)$ on conditional statement judgments across models. Llama 70B (vanilla) shows the strongest sensitivity, with a slope of (0.652 378 80.652\,378\,8 per unit increase in $P(B\mid A)$, suggesting that, with the absence of relevance influence, it is particularly sensitive to conditional probability. Out of the vanilla models, Qwen 72B shows the weakest effects (0.210 975 80.210\,975\,8).
How Conditional Probability and Relation Types Interact
For all models except Llama 70B (vanilla), relation type significantly modulates the effect of conditional probability, as indicated by highly significant interaction effects in the ANOVA results ($p<0.001$ for the vanilla models; see Tables 9 to 12 in the Appendix). Table 2 presents pairwise comparisons of interaction slopes, which indicate how much more strongly judgments increase with conditional probability for one relation type compared to another. These comparisons reveal distinct patterns across models. Llama 70B (vanilla) shows no significant slope differences (e.g., IRR $\times$ POS: $p=$ $0.6036$; others: $p=$ $1.0000$), reinforcing its insensitivity to relevance distinctions. Among the remaining models, two distinct trends emerge. In the smaller models (Llama 8B and Qwen 7B, both vanilla), positive relations yield the steepest slopes (e.g., Llama 8B slope differences: IRR $\times$ POS: 0.277-0.277; NEG $\times$ POS: 0.176-0.176; $p\leq 0.001$ respectively), suggesting a stronger probability effect for supportive conditionals. By contrast, larger models and all few-shot or CoT variants exhibit steeper slopes for negatively related conditionals than for positive ones (e.g., Llama 70B (few-shot): $0.195$, $p<0.001$) or, as in Qwen 72B (vanilla), show no significant difference between the two ($0.0295$, $p=$ $0.4197$). Not only does relevance modulate the effect of conditional probability, but the reverse also seems true: the influence of relation type depends on the underlying probability of the conditional statement. Pairwise comparisons of estimated marginal means at low, medium, and high probability levels (see Table 20 in the Appendix) reveal that in both Qwen models (7B and 72B, vanilla), all relation types differ significantly; except at low probability, where the irrelevant and negative conditions do not differ (e.g., Qwen 72B: $p=$ $0.2323$). A similar pattern appears in Llama 8B (vanilla), which shows no significant difference between negative and positive relations at low probability ($p=$ $0.1480$). These results suggest that while these LLMs reliably distinguish relation types at medium and high conditional probabilities, their sensitivity decreases at low probabilities; perhaps because relevance becomes harder to evaluate when the conditional itself seems implausible.
@@
Impact of Judgment Type
@@
Up to this point, we have analyzed if-probability and acceptability judgments of LLMs together. However, ANOVA results (see Tables 10 to 18 in the Appendix) suggest that some LLMs may differentiate between these two metrics. Only Llama 70B (vanilla) and Qwen 72B (vanilla) show no significant main effect of judgment type, though both approach the threshold of significance (Llama $p=$ 0.057 670.057\,67 Table 10; Qwen $p=$ 0.068 252 30.068\,252\,3 Table 12). Fixed effects estimates (using acceptability as the reference level, see Table 1) reveal no consistent trend among the vanilla models: Qwen 72B and Llama 8B give slightly lower ratings to if-probability (e.g., Qwen: 0.025 783 61-0.025\,783\,61), while Qwen 7B and Llama 70B show the opposite pattern (both: $0.05$). In contrast, all few-shot and CoT variants assign higher ratings to if-probability than to acceptability (see Table 1). This effect is most pronounced in Llama 8B (few-shot: 0.120 725 20.120\,725\,2) and smallest in Qwen 7B (few-shot: 0.054 886 550.054\,886\,55). These results suggest that prompting examples consistently shift model ratings upward for probability judgments relative to acceptability. However, for zero-shot, this distinction appears less stable and may depend on model size or architecture. Larger models like Llama 70B and Qwen 72B appear less influenced by judgment type overall, whereas prompting reintroduces or amplifies this sensitivity.
@@
All LLMs appear highly susceptible to few-shot and CoT prompting, though the direction of these effects varies across models. While few-shot prompting aids Llama 70B (vanilla) in incorporating relation types in its judgments (ANOVA: $p=$ 0.060 180.060\,18 to <0.0001<0.0001, see Tables 10 and 14 in the Appendix), it seems to introduce a negativity bias in other models: As can be seen in Table 2, all few-shot variants produce steeper interaction slopes for negatively related conditionals than for positively related ones (e.g., Llama 70B (few-shot): $0.195$ <0.001<0.001), or, as in Llama 8B (few-shot), eliminate the distinction entirely: $p=$ $0.5725$. Prompting effects are also evident in the pairwise comparisons of estimated marginal means across conditional probability levels (see Table 20 in the Appendix). In the vanilla Qwen models, relation type distinctions collapse only at low probability between irrelevant and negative conditionals. However, the few-shot and CoT variants additionally produce insignificant differences between negative and positive conditionals at high conditional probability (e.g., Qwen 72B CoT $p=$ $1.0000$), suggesting that the prompting examples may, in some cases, reduce sensitivity to fine-grained relevance contrasts. Notably, the Llama models do not exhibit this degradation under prompting, except for Llama 70B CoT, which converges negative and positive at medium conditional probability ($p=$ $0.4593$).
@@
4.1 Comparison Human vs. LLMs

### v2

@@ Abstract
Jasmin Orththanks: Equal contribution.1   Philipp Mondorf11footnotemark: 11, 2   Barbara Plank1, 2 1MaiNLP, Center for Information and Language Processing, LMU Munich, Germany 2Munich Center for Machine Learning (MCML), Munich, Germany jasmin.orth@campus.lmu.de, {p.mondorf, b.plank}@lmu.de
@@ 1 Introduction
While both sentences are grammatically sound, the second appears less acceptable: it is unlikely that Mark wearing socks has any effect on whether his TV is working. This notion of acceptability, or how appropriate a conditional statement seems in a given context, is often judged intuitively by humans (Douven and Verbrugge, 2012; Skovgaard-Olsen et al., 2016). Two main hypotheses aim to explain such intuitions. The conditional probability hypothesis holds that acceptability depends primarily on the conditional probability of $B$ given $A$ (Evans et al., 2003; Over et al., 2007). By this account, both conditionals above might be judged acceptable, as the probability of the TV being turned off given that the plug has been pulled, and the probability of the TV working given that Mark is wearing socks, are both high. In contrast, the evidentiality hypothesis emphasizes an additional influence of relevance or evidential connection encoded within the statement (Douven and Verbrugge, 2012; Skovgaard-Olsen et al., 2016). From this perspective, (i) is acceptable because pulling the plug is causally connected to the TV turning off, whereas (ii) is not, because wearing socks lacks such a connection. LLMs have demonstrated notable capabilities across diverse linguistic and cognitive tasks, including general reasoning (Hao et al., 2024; Mondorf and Plank, 2024a), linguistic inference (Clark et al., 2020), and causal reasoning (Dettki et al., 2025; Yu et al., 2024). However, despite increasing work on LLM reasoning, research on conditional acceptability judgments in LLMs remains virtually nonexistent. Most existing studies on LLMs and conditional statements focus on their causal, abductive, or inferential capabilities (Dettki et al., 2025; Holliday et al., 2024; Liu et al., 2023; Mondorf and Plank, 2024b), rather than on whether conditionals are judged contextually acceptable.
We address this gap by conducting a targeted investigation of conditional acceptability judgments in LLMs. Building on prior work from cognitive psychology (Skovgaard-Olsen et al., 2016), we assess the acceptability judgments of two model families: Llama 3.1 (8B and 70B) and Qwen 2.5 (7B and 72B), and examine how their judgments vary as a function of both the conditional probability and the evidential connection encoded within a statement. Our findings reveal that:
  • Acceptability judgments of models are influenced by both conditional probability and semantic relation, though the strength and consistency of this effect vary substantially by model family, size, and prompting technique.
@@ 2.1 Conditional Acceptability
The key concept of this study is conditional acceptability: the degree to which a conditional statement is perceived to be appropriate, plausible, or natural in context (Skovgaard-Olsen et al., 2016). While often associated with conditional probability $P(B\mid A)$, such as proposed by the conditional probability hypothesis (Evans et al., 2003; Oaksford and Chater, 2003; Over et al., 2007), the two can diverge significantly in practice. Take, for instance, the statement from the introduction: If Mark is wearing socks ($A$), then his TV will be working ($B$). This may be judged unacceptable or odd, even when both $P(A)$ and $P(B)$ are high, and even when $P(B\mid A)$ is high. The issue lies not in probability, but in the perceived irrelevance of $A$ to $B$. The statement violates intuitive notions of coherence or causality: wearing socks does not cause the TV to function, nor would not wearing them affect the outcome. As such, the distinction reflects the broader gap between correlation and causation (Barrowman, 2014). This disconnect has prompted extensive psychological research on how humans evaluate conditional statements (e.g., Douven, 2015; Evans and Over, 2004; Hardman and Macchi, 2003), suggesting that individuals often rely on a combination of semantic, contextual, and probabilistic cues when assessing conditionals (Berto and Özgün, 2021; Douven and Verbrugge, 2012; Over et al., 2007). Specifically, the evidentiality hypothesis argues that the semantic relation between antecedent $A$ and consequent BB—whether AA supports, contradicts, or is irrelevant to BB—plays a crucial role in human acceptability judgments (Douven and Verbrugge, 2012; Skovgaard-Olsen et al., 2016; Berto and Özgün, 2021). Returning to the earlier example, If Mark is wearing socks, then his TV will be working may be judged unacceptable not because $P(B\mid A)$ is low, but because there is no plausible explanatory or causal (supporting) link between the two clauses. In particular, Skovgaard-Olsen et al. (2016) find that humans generally judge conditionals with a supporting relation more acceptable than those with a contradicting or irrelevant relation.
@@ 2.2 Dataset
To investigate how LLMs assess the acceptability of conditional statements, we draw on a dataset from  Skovgaard-Olsen et al. (2016). Their study aims to understand how humans judge conditional acceptability using 144 conditional statements embedded in 12 short, realistic contexts. [^1] More information about the dataset and demographics can be found in Appendix A.1. Each context is tied to a unique everyday scenario, such as going on a date or attending a meeting.
To enable systematic analyses, the 12 conditionals that share the same context are organized along two dimensions. First, each conditional can be classified into one of three relation types, indicating the semantic connection between antecedent $A$ and consequent $B$: a positive relation ($POS$) means $A$ supports $B$, a negative relation ($NEG$) denotes that $A$ undermines or contradicts $B$, and an irrelevant relation ($IRR$) means $A$ is unrelated to $B$ (see Figure 2). Second, the conditionals vary in their prior probability: each scenario includes four levels based on combinations of high or low probability for both antecedent $P(A)$ and consequent $P(B)$ (e.g., high-high, high-low, low-high, low-low; see Appendix A.1 for details). This yields 12 unique configurations per context (3 relation types $\times$ 4 prior probabilities, as shown in Table 5 in the Appendix), resulting in 144 conditionals in total. The following example illustrates how relation types can be defined within a given context:
@@
As shown in Figure 2, a positive relation is illustrated by the conditional statement: If Nicole sunbathes on the beach, then she will get sunburned. Here, under the contextual constraint of her having sensitive skin, Nicole sunbathing actively causes her to get sunburned. In the negative example, the antecedent Nicole sunbathing on the beach contradicts her staying pale, while the irrelevant example pairs two unrelated ideas: knowing what a computer is has no causal connection to getting sunburned. These structured scenarios are designed to help participants apply real-world knowledge while simultaneously constraining interpretation, anchoring relations and probability judgments to a shared context. In the present study, we use this dataset to evaluate how LLMs judge the acceptability of conditionals and whether they exhibit similar patterns as humans. By prompting models with these same context-rich scenarios and conditionals, we can systematically assess their responses across relation types and probability levels.
@@ 3.1 Experimental Setup
We follow the approach of Skovgaard-Olsen et al. (2016) and present the 144 conditional statements described in Section 2.2 to LLMs in three distinct judgment tasks: (i) estimating the conditional probability $P(B\mid A)$ of the consequent given the antecedent, (ii.a) predicting the overall probability of the conditional $P\left(\text{``\emph{If A, then B}''}\right)$, and (ii.b) judging the acceptability of the conditional statement “If A, then B”. While tasks (ii.a) and (ii.b) present the LLMs with full conditional statements (e.g., ”If Mark is wearing socks, then his TV will be working”), the conditional probability task (i) presents the statement in a split format, consisting of a supposition such as “Suppose Mark is wearing socks” and an outcome, such as ”His TV will be working”, which needs to be judged (see Figure 3 lower bounding box).
@@
To analyze whether LLMs’ acceptability judgments align with a probabilistic or relevance-based interpretation, these tasks are designed to separate probabilistic reasoning from pragmatic or communicative considerations. The conditional probability task described above provides a reference point reflecting pure probabilistic reasoning: if large language models rely solely on conditional probability when judging conditional statements, their if-probability [^2] To make the distinction between the conditional probability $P(B\mid A)$ and the probability of the full conditional statement as clear as possible, we denote the probability of the full conditional statement the if-probability in this study. and acceptability ratings would be expected to closely match their corresponding $P(B\mid A)$ values. The split between acceptability and if-probability judgments further allows to determine whether assessments of conditional statements differ depending on the type of judgment. All tasks instruct the model to respond with a single number between 01000-100, indicating the degree of probability or acceptability. Full prompt templates are provided in Figures 8 to 10 in Appendix A.3.
Figure 4: Distribution of center-scaled judgments across humans and LLMs. Top row: overall distributions. Bottom row: distributions divided by relation type. Left: Llama 70B (vanilla), center: Qwen 72B (vanilla), right: human data obtained from the study by Skovgaard-Olsen et al. (2016).
Post-Processing
@@ 3.2 Prompts & Models
We aim to examine the impact of model family, model size, and prompting technique on the conditional acceptability judgments of LLMs. Therefore, we assess a total of four language models: Llama-3.1-8B-Instruct, [^3] https://huggingface.co/meta-llama/models Llama-3.1-70B- Instruct,footnotemark: Qwen2.5-7B-Instruct, [^4] https://huggingface.co/Qwen/collections and Qwen 2.5-72B-Instruct,footnotemark: using publicly accessible weights obtained from the Hugging Face platform. All models are prompted using a vanilla (zero-shot) and three-shot format, while the two larger models (Llama 70B and Qwen 72B) are also tested with chain-of-thought (CoT) prompting (Wei et al., 2022). For further details on models and prompts, please refer to Appendix A.2. Our code is publicly available at: https://github.com/mainlp/conditional-acceptability.
Cond. Prob. If-Prob. Acceptability
Vanilla
Llama 8B 0.7685 0.7724 0.7529
Llama 70B 0.9468 0.9344 0.9201
Qwen 7B 0.9768 0.9716 0.9450
Qwen 72B 0.9693 0.9673 0.9784
Few-shot
Llama 8B 0.7930 0.7538 0.7596
Llama 70B 0.9438 0.9556 0.8820
Qwen 7B 0.9542 0.9556 0.9596
Qwen 72B 0.9559 0.9437 0.9768
CoT
Llama 70B 0.9014 0.8497 0.8757
Qwen 72B 0.9331 0.9127 0.9661
Table 1: Intraclass correlation scores across five model runs for each judgment type, model, and prompt type.
Consistency of Model Responses
We generate model responses via nucleus sampling ($\text{top-p}=1.0$) and a temperature of $T=1.0$. As this decoding procedure is non-deterministic, we ensure the consistency of the models’ numerical ratings through repeated trials and a post-hoc correlation analysis. Each model is prompted five times per data sample to account for potential instabilities in ratings. By computing the intraclass correlation of these five ratings per sample, we further evaluate the reliability of the models’ judgments. As shown in Table 1, all models demonstrate good-to-excellent consistency across all judgment types and prompting strategies.
@@ 3.3 Statistical Methods
Mixed-effects models allow for the analysis of how conditional probability and relation type influence conditional acceptability judgments while accounting for additional sources of variation. In particular, different scenarios, participants, [^5] While we do not collect human data ourselves, we use human data from Skovgaard-Olsen et al. (2016) to compare LLM trends. and LLM sampling iterations introduce underlying variation that might otherwise obscure meaningful effects. Linear mixed-effects models account for this structure by incorporating both fixed effects (the predictors of interest) and random effects (uncontrolled variation across groups) (Brown, 2021). For further information on model fitting, please refer to Appendix A.5.
@@ 4 Results
Impact of Conditional Probability
For all LLMs, judgments of conditional statements are strongly influenced by the underlying conditional probability $P(B\mid A)$. ANOVA results (Tables 10 to 18 in the Appendix) confirm a highly significant main effect of conditional probability across all models ($p<0.0001$). As shown in Figure 6, judgments increase with rising conditional probability, a pattern further supported by fixed effects estimates in Table 2, which quantify the effect of $P(B\mid A)$ on conditional statement judgments across models. Llama 70B (vanilla) shows the strongest sensitivity, with a slope of 0.652 378 80.652\,378\,8 per unit increase in $P(B\mid A)$, suggesting that, with the absence of relevance influence, it is particularly sensitive to conditional probability. Out of the vanilla models, Qwen 72B shows the weakest effect (0.210 975 80.210\,975\,8).
@@
How Conditional Probability and Relation Types Interact
For all models except Llama 70B (vanilla), relation type significantly modulates the effect of conditional probability, as indicated by highly significant interaction effects in the ANOVA results ($p<0.001$ for the vanilla models; see Tables 9 to 12 in the Appendix). Table 3 presents pairwise comparisons of interaction slopes, which indicate how much more strongly judgments increase with conditional probability for one relation type compared to another. These comparisons reveal distinct patterns across models. Llama 70B (vanilla) shows no significant slope differences (IRR $\times$ POS: $p=$ $0.6036$; others: $p=$ $1.0000$), reinforcing its insensitivity to relevance distinctions. Among the remaining models, two distinct trends emerge. In the smaller models (Llama 8B and Qwen 7B, both vanilla), positive relations yield the steepest slopes (e.g., Llama 8B slope differences: IRR $\times$ POS: 0.277-0.277; NEG $\times$ POS: 0.176-0.176; $p\leq 0.001$ respectively), suggesting a stronger probability effect for supportive conditionals. By contrast, larger models and all few-shot or CoT variants exhibit steeper slopes for negatively related conditionals than for positive ones (e.g., Llama 70B (few-shot): $0.195$, $p<0.001$) or, as in Qwen 72B (vanilla), show no significant difference between the two ($0.0295$, $p=$ $0.4197$). Not only does relevance modulate the effect of conditional probability, but the reverse also seems true: the influence of relation type depends on the underlying probability of the conditional statement. Pairwise comparisons of estimated marginal means at low, medium, and high probability levels (see Table 20 in the Appendix) reveal that in both Qwen models (7B and 72B, vanilla), all relation types differ significantly; except at low probability, where the irrelevant and negative conditions do not differ (e.g., Qwen 72B: $p=$ $0.2323$). A similar pattern appears in Llama 8B (vanilla), which shows no significant difference between negative and positive relations at low probability ($p=$ $0.1480$). These results suggest that while these LLMs reliably distinguish relation types at medium and high conditional probabilities, their sensitivity decreases at low probabilities; perhaps because relevance becomes harder to evaluate when the conditional itself seems implausible.
Impact of Judgment Type
Up to this point, we have analyzed the if-probability and acceptability judgments of LLMs together. However, ANOVA results (see Tables 10 to 18 in the Appendix) suggest that some LLMs may differentiate between these two metrics. Only Llama 70B (vanilla) and Qwen 72B (vanilla) show no significant main effect of judgment type, though both approach the threshold of significance (Llama $p=$ 0.057 670.057\,67 Table 10; Qwen $p=$ 0.068 252 30.068\,252\,3 Table 12). Fixed effects estimates (using acceptability as the reference level, see Table 2) reveal no consistent trend among the vanilla models: Qwen 72B and Llama 8B give slightly lower ratings to if-probability (e.g., Qwen: 0.025 783 61-0.025\,783\,61), while Qwen 7B and Llama 70B show the opposite pattern (both: $0.05$). In contrast, all few-shot and CoT variants assign higher ratings to if-probability than to acceptability (see Table 2). This effect is most pronounced in Llama 8B (few-shot: 0.120 725 20.120\,725\,2) and smallest in Qwen 7B (few-shot: 0.054 886 550.054\,886\,55). These results suggest that prompting examples consistently shift model ratings upward for probability judgments relative to acceptability. However, for zero-shot, this distinction appears less stable and may depend on model size or architecture. Larger models like Llama 70B and Qwen 72B appear less influenced by judgment type overall, whereas prompting reintroduces or amplifies this sensitivity.
@@
All LLMs appear highly susceptible to few-shot and CoT prompting, though the direction of these effects varies across models. While few-shot prompting aids Llama 70B in incorporating relation types in its judgments (ANOVA: $p=$ 0.060 180.060\,18 to <0.0001<0.0001, see Tables 10 and 14 in the Appendix), it seems to introduce a negativity bias in other models: As can be seen in Table 3, all few-shot variants produce steeper interaction slopes for negatively related conditionals than for positively related ones (e.g., Llama 70B (few-shot): $0.195$, $p<0.001$), or, as in Llama 8B (few-shot), eliminate the distinction entirely: $p=$ $0.5725$. Prompting effects are also evident in the pairwise comparisons of estimated marginal means across conditional probability levels (see Table 20 in the Appendix). In the vanilla Qwen models, relation type distinctions collapse only at low probability between irrelevant and negative conditionals. However, the few-shot and CoT variants additionally produce insignificant differences between negative and positive conditionals at high conditional probability (e.g., Qwen 72B CoT $p=$ $1.0000$), suggesting that the prompting examples may, in some cases, reduce sensitivity to fine-grained relevance contrasts. Contrary to expectations, prompting examples do not appear to increase alignment with human behavior, which typically shows great sensitivity to relevance cues. Notably, the Llama models do not exhibit this degradation under prompting, except for Llama 70B CoT, which converges negative and positive at medium conditional probability ($p=$ $0.4593$).
@@
4.1 Comparison Humans vs. LLMs